Solve the given problem related to compound interest.
Find the balance if $6500 is invested at an annual interest rate of 6.5%, compounded annually, after the following number of years. (Round your answers to the nearest cent.)
(a) 5 years
$
(b) 12 years
$

Answer by lwsshak3(11628)

Solve the given problem related to compound interest.
Find the balance if $6500 is invested at an annual interest rate of 6.5%, compounded annually, after the following number of years. (Round your answers to the nearest cent.)
(a) 5 years
$
(b) 12 years
Compound interest formula: A=P(1+r/n)^nt, r=interest rate, n=number of compounding periods per year, A=amt after t-years
For 5 yrs
A=6500(1+.065/1)^1*
A=6500(1+.065)^5=6500*1.065^5=$8905.56
For 12 yrs
A=6500*1.065^12=$13,839.13